From a sack of fruit containing 2 apples, 3 oranges, and 2 bananas, a random sample of 4 pieces of fruit is selected. SupposePart b as well please
Answer #1

atples oamges bananas. Total Treoe is Sclcated 4 pice KASKCam apples x = 0, , 2, NO NO. o) P(x 0 O Lemnny Craung, 2 2 appl 22 a pie . P(x 2,) bamana 1 X Oyango, 1 harnana 2et a ppic 39 O banana UYounge appl 24 2 2 Magina! 3 1 X 12/35 2 1 1Aong Imal(x 3 Plx elx = O. (x 8, 2) x.+=2) (R2, 3 3 3 5 +y= 3) = tP(x,2) (R= 2,7=). 20 35 35 30 3 35 35 30 26 £3)